Ts planning meeting
This presentation is the property of its rightful owner.
Sponsored Links
1 / 11

TS planning meeting PowerPoint PPT Presentation


  • 43 Views
  • Uploaded on
  • Presentation posted in: General

TS planning meeting. 7 to 11 Nov 2011. Technical stop. Wk 45: Monday 7 th to Friday 11 th November Ramp down magnets (SL): Mon 8:00 – 8:30 (latest) RP search (Arturo and Philippe): Mon 8:30 – 9:15 Access starts Mon at 9:15 No access during the following periods:

Download Presentation

TS planning meeting

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Ts planning meeting

TS planning meeting

7 to 11 Nov 2011


Technical stop

Technical stop

  • Wk 45: Monday 7th to Friday 11th November

  • Ramp down magnets (SL): Mon 8:00 – 8:30 (latest)

  • RP search (Arturo and Philippe): Mon 8:30 – 9:15

  • Access starts Mon at 9:15

  • No access during the following periods:

    • Tuesday 7:00 to 9:00: periodic maintenance lift (AS703)

    • Thursday 6:00 to 9:00: leak test on FB pipe. No access, exceptions to be agreed with TC

    • Whole Friday morning (tbc depending on TPC progress): GS-ASE will install and test new sirens in UX25 for the evacuation alarm


Ts planning meeting

CV

  • Primary cooling:

    • Intervention SF2. Should not impact ALICE

  • Detector cooling (see next slide):

    • HMPID: replacement liquid sensor

    • SSD-SDD: replacement SSD T sensor

    • SPD: replacement pump


Activity d tector cooling

ActivityDétectorcooling

S.ROUSSEE EN/CV/DC


Detector activities

Detector activities

  • TPC:

    • Replace FEC (2 shifts from 08-14 and from 14-20 from Mon to Thu)

    • Fix leak on heat screen loop

  • TPC laser:

    • Access to laser hut: about 2-3 hours to check out the laser system

  • MTK:

    • Low voltage maintenance ST1-2 – Need light FASS

    • New CROCUS cable ST4 (SPIE). VIC needed, waiting for more info

    • Check BUS-patches in ST2

    • Remove cable lying on L4 gangway

  • ZDC:

    • Replace card in rack (C-side) ?

  • TRD:

    • Check switches in I-/O-/C-side racks

    • Check LV connections in sectors 10-11-16-17 (C-side). Tighten connections.


Detector activities1

Detector activities

  • MTG:

    • MT12-OUT-RPC6: HV cable check and eventual replacement; for this activity one nacelle is needed

    • Gas: activity in CR5 and eventually at the distribution rack in the cavern side C, level 0 (PH-DT). There is not gas activity foreseen on the RPCs

    • Check possibility to modify SAA3 shielding (Thu AM)

    • EMC: It should be only a half day of effort (probably on Mon afternoon):

    • Check and possible replacement of 1 RCU board from EMCal super module

    • Check communication link (cable) between one TRU module and STU module. Repair cable if necessary

    • Verify possibility to displace fibers (should be done in winter shutdown)


Detector activities2

Detector activities

  • TOF: replace DC/DC converters (Tue)

    • Move 2 SMs (1 and 10)

  • ITS (see Claudio’s presentation):

    • SSD:

      • Replace FEROM AD module on side A racks (Tue)

      • Check connection T sensor on cooling plant

    • SDD: replacement board Med5 (tbc)

    • SPD: ultrasound cleaning

  • ACO:

    • Calibration test to the electronics systems (rack O14 on Level-4)

    • Check 2 ACORDE modules that are not working properly

  • T0: checking laser system and electronics


Watcher schedule

Watcher schedule


Other activities

Other activities

  • TE-VSC: inspection GIS setup (Tue)

  • BE-ABP: check hydrostatic leveling system

  • EN-HE:

    • Mount crane electronics (Mon AM) and dismount it on Fri (tbc)

    • Load test cavern cranes (Tue, PM)

    • Maintenance SX2 crane (wk 48)

  • EN-EL: inspection PHOS cooling plant (UPS)

  • GA-ASE: maintenance PX24 access system (faulty biometric scan). Access will not be stopped

  • Intervention on the GSM and phone systems foreseen for all CERN on Wed

  • VIP visit (France 5) on Nov 9th at 11:00


Tc p2tech

TC & P2tech

  • Inspection 5 TPC bottom sectors

    • Thu 14:00. Ton, Borge and Luciano S.

  • Removal old fire detection pipes (Philippe)

  • Identify gas spare pipes (Philippe)

  • ADC access (Didier)

  • Designer (Al) + TC: check L3 fences, DCal passage, SAA3 modification, TRD installation

  • TC: look at TRD scaffolding (needed for Christmas)


Ts planning meeting

AOB

  • TRD: delivery next module around 17 Nov

  • Dcal:

    • Delivery beams to SX2 wk 44 and load test wk in 46

  • TRD:

    • Counterweight test wk 47

    • Reinforcement of frame ongoing. Test wk 48 (tbc)


  • Login